A electron passes through a magnetic field at 60° to the field at a velocity of 3.5×10⁶m/s. What is the magnitude of the force acting on the electron having a 0.45T magnetic field?

1
Expert's answer
2022-04-29T18:46:41-0400
"F=qvB\sin\theta\\
=1.6*10^{-19}*3.5*10^6*0.45*\sin 60^\circ\\
=2.2*10^{-13}\:\rm N"